OpenStudy (anonymous):

What is the solution to the equation log (2) x - log (2) 4 = 5

OpenStudy (anonymous):

log(2)4/4=5?

OpenStudy (anonymous):

log(2)x/4=5 sorry

OpenStudy (anonymous):

one way to do this is to know that \(\log_2(4)=2\) since \(2^2=4\) that gives you \[\log_2(x)+2=5\implies \log_2(x)=3\] then rewrite in exponential form as \(x=2^3\)

OpenStudy (anonymous):

ok it is minus, sorry start with \[\log_2(x)-2=5\implies \log_2(x)=7\implies x=2^7\]

OpenStudy (anonymous):

what property allows log(2)(4)=2?

OpenStudy (anonymous):

when i type it into my calculator it eqauls 1.204

OpenStudy (anonymous):

Simplify: log 2(x/4)=5 2^5=x/4 32=x/4 x=128

OpenStudy (anonymous):

The equation is : \[\Large \log_2x-\log_24=5 \] So : \[\Large \log_2x=5+\log_24\] But : \[\Large \log_24=\log_22^2=2\log_22=2\times1=2\] So the equation becomes : \[\Large \log_2x=5+2=7\] So : \[\Large x=2^7\]

OpenStudy (anonymous):

i see thakns alot guys!

OpenStudy (anonymous):

better than my teachers lol!